A 2.00 g sample of an iron ore is dissolved in acid and then reduced to Fe²⁺. It was then titrated against 0.025 M KMnO₄ solution and used 47.5 mL to reach the end point. Calculate the %Fe₃O₄ (231.54 g/mol) of the ore sample.



1
Expert's answer
2022-01-17T17:06:05-0500

Solution:

Balanced chemical equation (1): Fe3O4 → 3Fe3+→ 3Fe2+

Balanced chemical equation (2): 5Fe2+ + MnO4 + 8H+ → 5Fe3+ + Mn2+ + 4H2O


Calculate the moles of MnO4:

KMnO4 ⇌ K+ + MnO4

Moles of KMnO4 = Moles of MnO4

Moles of KMnO4 = Molarity of KMnO4 × Volume of KMnO4 solution

Moles of MnO4 = Moles of KMnO4 = (0.025 M × 0.0475 L) = 0.0011875 mol MnO4


According to the equation (2):

1 mol of MnO4 reacts with 5 moles of Fe2+

Thus, 0.0011875 mol of MnO4 reacts with:

Moles of Fe2+ = (0.0011875 mol MnO4) × (5 mol Fe2+ / 1 mol MnO4) = 0.0059375 mol Fe2+


According to the equation (1):

1 mol of Fe3O4 produces 3 mol of Fe2+

Thus, X mol of Fe3O4 produces 0.0059375 mol of Fe2+

Moles of Fe3O4 = (0.0059375 mol Fe2+) × (1 mol Fe3O4 / 3 mol Fe2+) = 0.001979 mol Fe3O4


Calculate the mass of Fe3O4:

The molar mass of Fe3O4 is 231.54 g/mol

Hence,

Mass of Fe3O4 = (0.001979 mol Fe3O4) × (231.54 g Fe3O4 / 1 mol Fe3O4) = 0.458 g Fe3O4


Calculate the %Fe3O4:

%Fe3O4 = (Mass of Fe3O4 / Mass of sample) × 100%

%Fe3O4 = (0.458 g / 2.00 g) × 100% = 22.9%

%Fe3O4 = 22.9%


Answer: %Fe3O4 = 22.9%